Date: Thu, 3 Dec 1998 16:15:00 -0800 (PST)
To: 2mass
Subject: IPAC 2MASS WG Meeting #174 Minutes
Cc: chas, stiening, bgreen,
Mime-Version: 1.0
Content-Transfer-Encoding: 7bit
Content-MD5: EOmc4sEIT2gSMTcE5HqbrQ==

           IPAC 2MASS Working Group Meeting #174 Minutes

Attendees: R. Cutri, S. Van Dyk, T. Jarrett, D. Kirkpatrick, R. Hurt,
           R. Tam, S. Wheelock, T. Evans, H. McCallon, R. Beck, 
           L. Fullmer, B. Nelson, J. Fowler, J. White, J. Rho, G. Kopan


1.) Project Update

2.) 2MASS Sampler

3.) Northern H- and Ks-Array Problems

4.) More on Hurt 1

5.) New and Standing Action Items 


     More treats from Hawaii started the meeting, courtesy of J. White.

1.) Project Update

     The possibility was raised of replacing the northern H-band array after
the Science Team Meeting in Dec.  There are currently bad pixels in the frames,
leading to holes in the coadds.  The number of bad pixels is increasing with
each thermal cycle.  The good news would be better images; the bad news, of 
course, is that such a replacement would lead to lots of IPAC work, with three 
to four weeks spent with such analyses as band-to-band adjustments, artifact 
tuning, etc.     

     In a day or two, dino will be diverted to southern data, starting with 
981010s.  There may be subtle problems with southern data only discovered by
processing recent nights.  Closer scrutiny is required.  The darks from 981001s
will be used, as we move toward running canonicals.  The nightly masks must be
compared to previous canonical masks.

     The agenda for the Science Team Meeting is not quite ready, but several 
IPAC team members will be giving presentations.

2.) 2MASS Sampler

     The Sampler documentation contributions have all been submitted by 2MASS
team members.  These need to all be compiled, and perusal by Science Team
members is required.  R. Cutri asked yet again whether everyone has been busy
exercising CatScan and the Survey Visualizer.  Any comments on these two tools
should be sent to IRSA, being sure to cc R. Cutri.  A Sampler feedpack webpage
is being formulated, for Sampler users to fill out.  It needs to be developed
carefully, to obtain necessary information from users.

3.) Northern H- and Ks-Array Problems
     G. Kopan reported on depressions in northern Ks coadds, due to hot pixels
and erratic offsets.  A few more pixels will need to be masked off.  Discussion
of the H-band problems continued, namely holes 2 or 3 pixels a side, and the
question of the "absolute badness" of the array.  Some confusion in treating
problems with the H-array ensued, from 981023n and the substitution of the 
EPROM, when col 129 was turned off and then turned back on.  About 20 pixels
went bad after the warmup on 981018n.  (See Action Items below.)

     The question is, how are the bad pixels affecting the photometry?  R. Tam
is to look for missing aperture photometry, due to dead pixels, as a function 
of cross-scan.  Further analyses by G. Kopan, S. Wheelock, and L. Fullmer will 
investigate increased noise and false detections, and increased photometric

     Less than one-fourth of the sky will be covered by a better H-array in the
north, if a substitution takes place.  The risks involved are numerous, and
the IPAC team will have to learn all the idiosyncracies of the new array.

     J. Fowler continued the discussion of the holes.  He suggested that they
might be cured by resetting the thresholds to non-infinite values in DFLAT,
which would just require changing namelist parameters.  R. Cutri asked him to
write up a proposal for action on this.  Some test scans could be run with this
fix.  J. Fowler suggested that a whole night should be run, but R. Cutri 
reminded him that this would require a lot of time for analysis.  J. Fowler
inquired when this problem first appears.  G. Kopan indicated that it is
apparent by 981023n, but could be earlier, during the recent warmup period and
the change of electronics.
4.) More on Hurt 1

     R. Hurt, fresh from a Palomar 200-in run, obtained a deep red spectrum of
the possible new obscured galaxy.  The spectrum looks like a galaxy, and there
may be weak H-alpha.  New HI observations of the field show a velocity and
velocity dispersion consistent with a galaxy at this location, but the HI
redshift suggests it is not a Local Group member.
5.) New and Standing Action Items

%                                                      %
%                 ACTION ITEMS                         %
%                                                      %

      Action Item for J. Fowler, G. Kopan, and R. Cutri: to figure out masking
for various recent breakpoints.

      Action Item for all team members:  Have a good Thanksgiving holiday.

(Standing Action Items:)

a) Team members should continue testing the IRSA tools and the Sampler release.

b) Team members should continue to analyze data for Spring data release.

c) Team members are reminded that performance reports covering October and 
November will be due on Dec 2.